OpenStudy (anonymous):

write an equation of the line containing the given point and parallel to given line. Express your answer in the form of y=mx+b (7,8) x+3y=4

OpenStudy (kropot72):

The first step is to rearrange the given equation into the form y = mx + b where m is the slope. This value of slope must be the same in the equation you need to find. x + 3y = 4 Moving x to the right hand side gives: 3y = -x + 4 .............(1) Can you now put equation (1) into the form y = mx + b?

OpenStudy (anonymous):

-x/3+4/3

OpenStudy (kropot72):

Yes the given equation now becomes: y = -x/3 + 4/3 where the slope m = -1/3 So the equation that you need to find is now y = -x/3 + b ..............(2) and you must find the value of b by substituting the coordinates of the given point (7, 8) into equation (2). Can you do that?

OpenStudy (anonymous):

31/3=b?

OpenStudy (kropot72):

Yes, 31/3 or 10 1/3 is correct.

OpenStudy (anonymous):

i dont remember how i got the 31/3 to =b

OpenStudy (kropot72):

@tishagwen You need to substitute x = 7 and y = 8 into the equation y = -x/3 + b and solve for b.

OpenStudy (anonymous):

so the slope does or doesn't get included

OpenStudy (anonymous):

8=-7/3+b?

OpenStudy (anonymous):

do i get b by itself

OpenStudy (kropot72):

Correct so far. Now add 7/3 to both sides.

OpenStudy (anonymous):

-8/3

OpenStudy (anonymous):

wow.. nevermind.. hit wrong number in calc.. i got the 31/3 again

OpenStudy (kropot72):

8 + 7/3 = b -7/3 + 7/3 = b b = 8 + 7/3 = 24/3 + 7/3 = 31/3

OpenStudy (anonymous):

thank you.. my internet disconnected the other day and i hadnt had a chance to get back on here and review it.. and i couldnt remember how i got 31/3

OpenStudy (kropot72):

You're welcome :)